Part B Classify the bonds formed between each pair of elements based on electronegativity values in the following table: Electronegativity 1.3 1.5 2.1 2.1 2.5 2.5 2.8 3.0 3.5 4.0 Drag the appropriate items to their respective bins. Element Se Be H P S I B C O F View Available Hint(s) Reset Help H and Br H and P S and S and o Sc and o Be and F Be and Cl Polar covalent Nonpolar covalent onic

Part B Classify the bonds formed between each pair of elements based on electronegativity values in the following table: 


Drag the appropriate items to their respective bins. 
Answer #1

Bond type. 1 1 s-0 Bond Electronegativity difference 1201 – 2.81 = H-p 0.7 Polar covalent 1 2.) - 2011 = Non polar covalent S